OpenStudy (sleepyjess):

Just need someone to check my answer please Simplify (1 − cos x)(1 + cos x). 1 cos^2 x <------ sin^2 x tan^2 x

jimthompson5910 (jim_thompson5910):

that is incorrect

OpenStudy (anonymous):

\(1-\cos^{2}x\)

jimthompson5910 (jim_thompson5910):

when you used the difference of squares formula, what did you get? (1 − cos x)(1 + cos x) = ( )^2 - ( )^2 what goes in each parenthesis?

OpenStudy (anonymous):

then use the pythagorean identity \(\cos^{2}x + \sin^{2}x = 1\)

OpenStudy (sleepyjess):

it is 1 isn't it? because it is cos x^2 -cosx^2

hero (hero):

\(cos x^2 -cosx^2 = 0\)

OpenStudy (anonymous):

nope... and you may want to use the equation editor to write more clearly.

OpenStudy (sleepyjess):

so what is the answer and how do i get to it?